# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v3

@@ IV Thermodynamic Constraints and Physical Limits
This section establishes the physical constraints on information processing in knowledge systems, connecting our information-theoretic framework to thermodynamics via Landauer’s principle.
@@ IV-A Landauer’s Principle and Energy Bounds
Theorem 8 (Landauer Energy Bound for Computation).
@@
Under the assumption that each inference step constitutes an ideal logically irreversible operation, deriving formula $\phi$ from knowledge base $S_{O}$ requires minimum energy:
$E_{\text{compute}}(\varphi|S_{O})\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (72)
where $k_{B}=1.38\times 10^{-23}$ J/K is Boltzmann’s constant and $T$ is the operating temperature.
Each inference step in deriving $\varphi$ from $S_{O}$ performs an irreversible logical operation. By Landauer’s principle [landauer1961irreversibility], each irreversible operation erasing or transforming 1 bit of information requires minimum energy $k_{B}T\ln(2)$.
Consider a derivation with $\text{Dd}(\varphi|S_{O})$ steps. Each step:
@@
    Applies an inference rule $r_{i}:\{\psi_{1},\ldots,\psi_{k}\}\vdash\psi^{\prime}$
@@
    Writes conclusion $\psi^{\prime}$ to working memory
  3. 3.
    Erases intermediate results (at least 1 bit per step)
The total bit erasures across all steps is at least $\text{Dd}(\varphi|S_{O})$. By Landauer’s principle, total energy:
$E_{\text{compute}}\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (73)
Theorem 9 (Storage Energy Bound).
Maintaining storage $S$ containing $|S|$ bits for time $t$ requires energy:
$E_{\text{storage}}(S,t)\geq|S|\cdot k_{B}T\ln(2)\cdot\frac{t}{t_{\text{refresh}}}$ (74)
where $t_{\text{refresh}}$ is the thermal decoherence time.
Physical storage media suffer thermal noise causing bit flips at rate $\lambda_{\text{flip}}\sim\exp(-E_{\text{barrier}}/k_{B}T)$. The refresh period is $t_{\text{refresh}}=1/\lambda_{\text{flip}}$. Each refresh cycle requires re-writing $|S|$ bits, each costing $k_{B}T\ln(2)$ energy. Over time $t$, number of refreshes is $\lceil t/t_{\text{refresh}}\rceil$, yielding the stated bound. ∎
IV-B Energy-Time-Space Triality Bound
We now prove our main physical constraint theorem.
Theorem 10 (Fundamental Triality Bound).
For any hybrid strategy $\Pi=(S,C,R)$ answering query distribution $P_{Q}$ over knowledge base $S_{O}$ in the information-rich regime:
$E(\Pi)\cdot T(\Pi)\cdot M^{-1}(\Pi)\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (75)
This bound is asymptotically tight and achieved by the optimal hybrid strategy in the information-rich regime (see Section VI).
[Preamble] Since queries are in the information-rich regime:
$H(Q|S_{O})\gg\log_{2}(|\text{Atom}(S_{O})|),\log_{2}(Dd(Q|S_{O}))$ (76)
By Corollary LABEL:cor:Asymptotic_Equivalence, this implies:
$Dd(Q|S_{O})\cdot\ln(2)\approx H(Q|S_{O})\quad[\text{asymptotically}]$ (77)
We analyze three cases using this approximation.
Case 1: Pure Computation Strategy ($\Pi_{c}$)
Setup: - No precomputed storage: $S=\emptyset$, hence $M(\Pi_{c})\to 0$.
- Time cost: Average number of inference steps:
$T(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[Dd(q|S_{O})]=Dd(Q|S_{O})$ (78)
- Energy cost: By Theorem 8,
$E(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[E_{\text{compute}}(q|S_{O})]\geq Dd(Q|S_{O})\cdot k_{B}T\ln(2)$ (79)
Energy-Time Product: Since we are in the information-rich regime, we use the asymptotic approximation:
$Dd(Q|S_{O})\cdot\ln(2)\approx H(Q|S_{O})$ (80)
Thus,
$\begin{split}E(\Pi_{c})\cdot T(\Pi_{c})&\geq(Dd(Q|S_{O}))^{2}\cdot k_{B}T\ln(2)\\ &\geq\Omega(H(Q|S_{O})^{2}/\ln(2))\cdot k_{B}T\ln(2)\\ &=\Omega(H(Q|S_{O})^{2}\cdot k_{B}T)\end{split}$ (81)
Case 2: Pure Storage Strategy ($\Pi_{s}$)
Setup: - Precompute all answers: $S=\{(q,A(q,S_{O})):q\in\text{supp}(P_{Q})\}$. - Storage requirement:
$M(\Pi_{s})=H(Q|S_{O})$ (82)
- Maintenance energy: By Theorem 9, over average query inter-arrival time $t_{\text{avg}}=1/\lambda_{Q}$:
$E(\Pi_{s})=H(Q|S_{O})\cdot k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (83)
Energy-Storage Product:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})=k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (84)
For reasonable systems where $t_{\text{avg}}\sim t_{\text{refresh}}$:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})\geq\Omega(k_{B}T\ln(2))$ (85)
In the information-rich regime, this bound holds as stated.
Case 3: Hybrid Strategy ($\Pi_{h}$)
Setup: - We consider a hybrid strategy that balances both precomputation and on-demand computation.
Analysis: By the Storage-Computation Duality Theorem (Theorem LABEL:thm:duality):
$\min[\text{Storage}]+\min[\text{Computation}]\geq H(Q|S_{O})\cdot(1+\ln(2))$ (86)
In the information-rich regime, the total cost can be approximated as:
$E(\Pi_{h})\cdot T(\Pi_{h})\cdot M^{-1}(\Pi_{h})\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (87)
This completes the analysis for all three cases, showing consistency with the triality bound.
Corollary 6 (Critical Storage Capacity).
The critical storage capacity is:
$M_{\text{critical}}=\frac{H(Q)}{\log\left(\frac{E_{\text{budget}}}{H(Q|S_{O})\cdot k_{B}T\ln(2)}\right)}$ (88)
Below this capacity, storage is the bottleneck; above it, computation dominates.

### v4

@@ IV THERMODYNAMIC CONSTRAINTS AND PHYSICAL LIMITS
This section establishes the physical constraints on information processing in knowledge systems, connecting our information-theoretic framework to thermodynamics via Landauer’s principle and the physical cost assumptions of Section II.
@@ IV-A Landauer’s Principle and Energy Bounds
We now relate derivation depth and storage size to minimal energy requirements under the classical thermodynamic regime of Assumption 3.
Recall that, in this regime, each logically irreversible operation that commits or erases one bit of information dissipates at least $k_{B}T\ln 2$ joules of heat, where $k_{B}$ is Boltzmann’s constant and $T$ is the operating temperature. In Section III-A we fixed the per-step information cost to be $\lambda\;=\;\ln 2\ \text{nats}$, so that each inference step in a derivation is modeled as committing or erasing $\Theta(1)$ bit of information, and, in particular, we normalize units so that each step incurs a Landauer cost of at least $k_{B}T\ln 2$.
Proposition 4 (Landauer energy bound for computation).
@@
Under the assumption that each inference step constitutes a logically irreversible operation in the classical thermodynamic regime (Assumption 3), deriving a formula $\varphi$ from a knowledge base $S_{O}$ requires at least the following energy:
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;Dd(\varphi\mid S_{O})\,\cdot\,k_{B}T\ln 2,$ (98)
where $Dd(\varphi\mid S_{O})$ is the derivation depth of $\varphi$ from $S_{O}$ in the sense of Definition 6.
By Assumption 3 and the discussion above, we model each inference step as a logically irreversible operation that commits or erases at least one bit of information. By Landauer’s principle [landauer1961irreversibility], each such operation incurs an energy cost of at least $k_{B}T\ln 2$.
Consider a derivation of $\varphi$ from $S_{O}$ of minimal derivation depth $Dd(\varphi\mid S_{O})$. This derivation consists of exactly $Dd(\varphi\mid S_{O})$ inference steps. Since each step is modeled as performing at least one bit’s worth of logically irreversible information processing, the total number of bit erasures (or committed bits) across the derivation is lower bounded by
$\text{bit-erasures}\;\geq\;Dd(\varphi\mid S_{O}).$
By Landauer’s principle, the total energy dissipated by these operations thus satisfies
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;Dd(\varphi\mid S_{O})\,k_{B}T\ln 2,$
which is exactly (98). ∎
In terms of the derivation entropy defined in Section III-A, we immediately obtain a compact form.
Remark 9 (Derivation entropy as thermodynamic cost).
Under Assumption 3 we have, by Corollary 3,
$H_{\mathrm{derive}}(\varphi\mid S_{O})\;=\;Dd(\varphi\mid S_{O})\,\ln 2\quad\text{(in nats)}.$
Combining this with Proposition 4 yields
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;k_{B}T\,H_{\mathrm{derive}}(\varphi\mid S_{O}),$ (99)
so that derivation entropy directly lower-bounds the minimal thermodynamic energy required to carry out the corresponding inference.
We now complement the dynamic computation cost with an energy bound for maintaining stored information over time.
Proposition 5 (Energy bound for maintaining storage).
In a simple Arrhenius model of thermally noisy storage, maintaining a stored state $S$ with $|S|$ effective bits over a time horizon $t$ requires at least
$E_{\mathrm{storage}}(S,t)\;\gtrsim\;|S|\,k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}},$ (100)
where $t_{\mathrm{refresh}}$ denotes the characteristic refresh interval implied by the thermal flip rate of the storage medium.
We adopt a coarse-grained Arrhenius model for thermal bit flips. Let $E_{\mathrm{barrier}}$ denote the energy barrier separating the two logical states of a physical bit cell. Classical transition-state theory then suggests that the bit-flip rate satisfies
$\gamma_{\mathrm{flip}}\;\approx\;\gamma_{0}\exp\!\Bigl(-\frac{E_{\mathrm{barrier}}}{k_{B}T}\Bigr),$
for some attempt frequency $\gamma_{0}$ [atkins2014physical]. The characteristic time between spontaneous flips is thus
$t_{\mathrm{refresh}}\;\approx\;\frac{1}{\gamma_{\mathrm{flip}}},$
which we interpret as the required refresh interval to reliably maintain the stored logical state against thermal noise.
In a simple refresh model, each refresh cycle rewrites all $|S|$ bits in $S$ to their intended logical values. By Landauer’s principle, each rewritten bit incurs at least $k_{B}T\ln 2$ of energy dissipation, so a single refresh of $S$ requires at least
$E_{\mathrm{per\;refresh}}\;\geq\;|S|\,k_{B}T\ln 2.$
Over a total time horizon $t$, the number of such refresh cycles is on the order of $t/t_{\mathrm{refresh}}$. Hence the total energy dissipated in maintaining $S$ satisfies
$E_{\mathrm{storage}}(S,t)\;\gtrsim\;\frac{t}{t_{\mathrm{refresh}}}\,|S|\,k_{B}T\ln 2,$
which is (100). Here the symbol $\gtrsim$ emphasizes that we are working with a lower bound under simplifying assumptions: (i) independent bit flips with Arrhenius rate, and (ii) full rewriting of all $|S|$ bits at each refresh cycle. More sophisticated coding or hardware techniques may improve constant factors but cannot avoid the linear dependence on $|S|$ and $t/t_{\mathrm{refresh}}$ in this regime. ∎
Remark 10 (Link to amortized storage cost).
The storage size $|S|$ already appears in the information-theoretic amortized-cost decomposition of Section III-C, where the expected per-query information cost is
$E[\mathrm{Cost}]\;=\;\frac{|S|}{N}\;+\;E_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
for a workload of $N$ queries drawn from $P_{Q}$. Proposition 5 shows that this bit-length $|S|$ also controls a minimal thermodynamic cost for maintaining the corresponding physical memory:
$\frac{E_{\mathrm{storage}}(S,t)}{N}\;\gtrsim\;\frac{|S|}{N}\,k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}},$
which is directly analogous to the $\frac{|S|}{N}$ term above. Together with (99), this yields a chain of correspondences
$\text{derivation depth }Dd(\cdot)$ $\;\longrightarrow\;H_{\mathrm{derive}}(\cdot)\ \text{(nats)}$
$\;\longrightarrow\;E_{\mathrm{compute}}(\cdot)\ \text{(J)},$
and
$\text{storage size }|S|\ \text{(bits)}$ $\;\longrightarrow\;\frac{|S|}{N}\ \text{(bits/query)}$
$\;\longrightarrow\;\frac{E_{\mathrm{storage}}(S,t)}{N}\ \text{(J/query)}.$
Thus, the storage–computation tradeoffs derived in Section III not only constrain algorithmic and information-theoretic costs, but also translate into minimal thermodynamic costs in physically implemented knowledge systems.
IV-B Storage-Computation Energy Tradeoff
Having established the information-theoretic foundations of mixed strategies in Section III, we now derive fundamental lower bounds on the total energy cost of semantic information processing. These results connect the abstract information measures to physical thermodynamic constraints.
We consider a knowledge system operating over a time horizon during which $N$ queries are processed. Let $P_{Q}$ be a computable probability distribution over queries, and let $Q$ denote a random query drawn from $P_{Q}$. We assume the system operates in the classical thermodynamic regime (Assumption 3).
Definition 11 (Amortized Energy Cost).
For a mixed strategy $\Pi=(S,C,R)$ processing $N$ queries over time $t$, the amortized energy cost per query is defined as:
$\bar{E}(\Pi):=\frac{E_{\mathrm{storage}}(S,t)}{N}+\mathbb{E}_{q\sim P_{Q}}\left[E_{\mathrm{compute}}(q|S_{O}\cup S)\right]$ (101)
where $E_{\mathrm{storage}}(S,t)$ is the total energy required to maintain storage $S$ for time $t$, and $E_{\mathrm{compute}}(q|S_{O}\cup S)$ is the energy required to derive the answer to process query $q$ from the augmented knowledge base $S_{O}\cup S$.
Throughout this section, we assume the system operates in the standard refresh regime, meaning that the time horizon $t$ and query count $N$ satisfy $t/(N\cdot t_{refresh})=\Theta(1)$, i.e., the storage is refreshed a constant number of times per query on average.
Definition 12 (Normalized Information Cost).
For a mixed strategy $\Pi=(S,C,R)$, the normalized information cost per query is:
$\bar{I}(\Pi):=\frac{|S|}{N}+\mathbb{E}_{q\sim P_{Q}}\left[H_{\mathrm{derive}}(q|S_{O}\cup S)\right]$ (102)
where $|S|$ denotes the size of auxiliary storage in bits, and $H_{\mathrm{derive}}(q|S_{O}\cup S)=D_{d}(q|S_{O}\cup S)\cdot\ln 2$ is the derivation entropy.
The normalized information cost $\bar{I}(\Pi)$ serves as an information-theoretic proxy for energy consumption, with the conversion to physical energy units given by Propositions 4 and 5.
We now state and prove the main theorem of this section, which establishes a fundamental lower bound on the resources required by any mixed strategy.
Assumption 4 (Uniform Polynomial-Depth Regime).
Let $P_{Q}$ be a query distribution over a query set $\mathcal{Q}$ with respect to knowledge base $S_{O}$. We say that $P_{Q}$ satisfies the uniform polynomial-depth regime with respect to a mixed strategy $\Pi=(S,C,R)$ if there exists a polynomial $p(\cdot)$ such that
$D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S):=\\ \sup_{q\in\mathrm{supp}(P_{Q})}D_{d}(q\mid S_{O}\cup S)\leq p\bigl(|\mathrm{Atom}(S_{O}\cup S)|\bigr).$ (103)
In particular, writing $m_{S}:=|\mathrm{Atom}(S_{O}\cup S)|$, we have $\log D_{d}^{\sup}=O(\log m_{S})$ and hence
$L_{S}:=\log(m_{S}+D_{d}^{\sup})=\Theta(\log m_{S}).$ (104)
Remark 11 (Scope of Assumption 4).
The uniform polynomial-depth regime is satisfied by most practical query classes, including:
  • Datalog queries with bounded recursion depth [ceri1989you];
  • SPARQL queries over RDF graphs with bounded property path length [perez2009semantics];
  • Template-based natural language queries compiled to first-order logic;
  • Bounded-depth inference in description logics (e.g., $\mathcal{ALC}$, $\mathcal{SHIQ}$).
The key requirement is that the inference depth does not grow super-polynomially with the size of the knowledge base—a condition that holds for essentially all tractable reasoning tasks. More precisely, for these query classes, the derivation depth is bounded by a fixed polynomial in the number of base facts, independent of the query complexity.
Theorem 6 (Storage-Computation Information Tradeoff).
Let $S_{O}$ be a knowledge base and let $P_{Q}$ be a computable query distribution satisfying the information-rich regime (Definition 8). Let $\Pi=(S,C,R)$ be any mixed strategy satisfying assumptions (SA1)–(SA3) and the uniform polynomial-depth regime (Assumption 4).
Define the effective complexity parameter:
$L_{S}:=\log\Bigl(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S)\Bigr),$ (105)
where $D_{d}^{\sup}(P_{Q}|S_{O}\cup S)$ is as defined in Assumption 4.
If the total number of queries satisfies $N=\Omega(L_{S})$, then there exists a constant $c>0$ depending only on the proof system and encoding scheme such that:
$\bar{I}(\Pi)\geq c\cdot\frac{H(Q\mid S_{O})}{L_{S}}-O(1),$ (106)
where $H(Q\mid S_{O})$ denotes the conditional Shannon entropy of $Q$ given the knowledge base $S_{O}$.
The proof proceeds in five steps.
Step 1: Lower bound on derivation entropy. By the lower bound in Theorem 4, for any query $q\in\text{supp}(P_{Q})$ in the information-rich regime, the derivation entropy satisfies:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)=D_{d}(q\mid S_{O}\cup S)\cdot\ln 2\\ \geq c_{1}\frac{H(q\mid S_{O}\cup S)}{L(q)}-O(1),$ (107)
where $c_{1}>0$ is the constant from Theorem 4, $H(q\mid S_{O}\cup S):=K(q\mid\langle\mathrm{Atom}(S_{O}\cup S)\rangle)\cdot\ln 2$ is the conditional algorithmic entropy, and
$L(q):=\log\bigl(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}(q\mid S_{O}\cup S)\bigr).$ (108)
Step 2: Uniformization via the polynomial-depth assumption. Under Assumption 4, we have for all $q\in\mathrm{supp}(P_{Q})$:
$D_{d}(q\mid S_{O}\cup S)\leq D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S),$ (109)
and consequently,
$L(q)\leq L_{S}\quad\text{for all }q\in\mathrm{supp}(P_{Q}).$ (110)
Moreover, since both $|\mathrm{Atom}(S_{O}\cup S)|$ and $D_{d}^{\sup}$ are polynomially related under Assumption 4, we have:
$L(q)=\Theta(L_{S})\quad\text{for all }q\in\mathrm{supp}(P_{Q}).$ (111)
Therefore, there exists a constant $c_{1}^{\prime}>0$ (absorbing the $\Theta(\cdot)$ factor into $c_{1}$) such that for all $q\in\mathrm{supp}(P_{Q})$:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\geq c_{1}^{\prime}\frac{H(q\mid S_{O}\cup S)}{L_{S}}-O(1).$ (112)
Taking expectations over $q\sim P_{Q}$:
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\\ \geq c_{1}^{\prime}\frac{\mathbb{E}_{q\sim P_{Q}}\bigl[H(q\mid S_{O}\cup S)\bigr]}{L_{S}}-O(1),$ (113)
where the $O(1)$ term remains $O(1)$ after expectation since it is uniform over all $q\in\mathrm{supp}(P_{Q})$.
Step 3: Relating Kolmogorov complexity to Shannon entropy. For a computable distribution $P_{Q}$, the expected Kolmogorov complexity relates to Shannon entropy via the classical correspondence (see, e.g., [LiVitanyi2019Kolmogorov, cover2006elements]):
$\mathbb{E}_{q\sim P_{Q}}\bigl[K(q\mid S_{O}\cup S)\bigr]\geq H(Q\mid S_{O},S)-O(1),$ (114)
where $H(Q\mid S_{O},S)$ is the conditional Shannon entropy of $Q$ given both $S_{O}$ and $S$, and the $O(1)$ term accounts for the description complexity of the distribution $P_{Q}$ (which is assumed computable and hence has finite Kolmogorov complexity).
By the chain rule for Shannon entropy:
$H(Q\mid S_{O},S)=H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (115)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between the auxiliary storage $S$ and the query $Q$ given $S_{O}$.
Combining with (113):
$\mathbb{E}\bigl[H_{\mathrm{derive}}(Q\mid S_{O}\cup S)\bigr]\\ \geq c_{1}^{\prime}\frac{H(Q\mid S_{O})-I(S;Q\mid S_{O})}{L_{S}}-O(1).$ (116)
Step 4: Bounding the mutual information. By assumption (SA1), the auxiliary storage $S$ is a valid logical encoding, which implies:
$I(S;Q\mid S_{O})\leq H(S\mid S_{O})\leq H(S)\leq|S|,$ (117)
where the first inequality is the standard property that mutual information is bounded by marginal entropy, the second follows from conditioning reducing entropy, and the third follows from the fact that Shannon entropy is bounded by the description length.
Step 5: Combining storage and computation costs. From the definition of $\bar{I}(\Pi)$ and inequalities (116) and (117):
$\bar{I}(\Pi)=\frac{|S|}{N}+\mathbb{E}[H_{\mathrm{derive}}(Q|S_{O}\cup S)]\\ \geq\frac{|S|}{N}+c_{1}^{\prime}\frac{H(Q|S_{O})-|S|}{L_{S}}-O(1).$ (118)
We analyze this expression by considering the extremal cases.
Case (a): Pure computation ($|S|=0$). In this case,
$\bar{I}(\Pi)\geq c_{1}^{\prime}\frac{H(Q|S_{O})}{L_{S}}-O(1).$
Case (b): Pure storage ($|S|=H(Q|S_{O})+O(1)$). By assumption (SA1), any useful storage satisfies $|S|\leq H(Q|S_{O})+O(\log|S_{O}|)$. Taking $|S|=H(Q|S_{O})$, we obtain
$\bar{I}(\Pi)\geq\frac{H(Q|S_{O})}{N}+0-O(1)=\frac{H(Q|S_{O})}{N}-O(1).$
Under $N=\Omega(L_{S})$, this yields $\bar{I}(\Pi)\geq\Omega(H(Q|S_{O})/L_{S})-O(1)$.
General case. For intermediate values of $|S|$, the right-hand side of (118) is a linear function of $|S|$ with slope $\frac{1}{N}-\frac{c_{1}^{\prime}}{L_{S}}$. Since $N=\Omega(L_{S})$, this slope is $O(1/L_{S})$ in magnitude. In either case (slope positive or negative), the minimum of the bound over $|S|\in[0,H(Q|S_{O})]$ is achieved at one of the endpoints, which we have already analyzed.
Combining Cases (a) and (b), there exists a constant $c>0$ such that
$\bar{I}(\Pi)\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}-O(1),$
completing the proof. ∎
Remark 12 (Role of the Supremum in $L_{S}$).
The use of $D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S)$ rather than, say, the expected value $\mathbb{E}[D_{d}(Q\mid S_{O}\cup S)]$ in the definition of $L_{S}$ serves two purposes:
@@ IV-B Storage-Computation Energy Tradeoff
    Uniformity: The lower bound (106) holds for all queries in the support of $P_{Q}$, not just on average. This is essential for worst-case guarantees.
@@
    Asymptotic equivalence: Under the polynomial-depth regime (Assumption 4), the logarithmic factor satisfies
    $\log\bigl(m+\mathbb{E}[D_{d}]\bigr)=\Theta\bigl(\log(m+D_{d}^{\sup})\bigr)=\Theta(\log m),$
    where $m=|\mathrm{Atom}(S_{O}\cup S)|$. Thus, the choice between supremum and expectation affects only the implicit constants, not the asymptotic behavior.
In practice, for query distributions concentrated on a polynomial-depth class, the distinction is immaterial.
Remark 13 (Tightness of the Bound).
The lower bound in Theorem 6 is tight up to constant factors. By Theorem 5, there exists a strategy $\Pi^{*}$ achieving:
$\bar{I}(\Pi^{*})\leq C\cdot\frac{H(Q|S_{O})}{L_{S^{*}}}+O(1)$ (119)
for some constant $C>0$, where $L_{S^{*}}$ is the value of $L_{S}$ corresponding to the optimal strategy $\Pi^{*}$. Thus, the optimal normalized information cost satisfies:
$\bar{I}^{*}=\Theta\left(\frac{H(Q|S_{O})}{L^{*}}\right)$ (120)
where $L^{*}=\Theta(L_{S})$ for the optimal strategy.
We now translate the information-theoretic bound into physical energy units using the thermodynamic correspondences established in Section IV-A.
Corollary 6 (Thermodynamic Energy Lower Bound).
Under the conditions of Theorem 6, additionally assuming the classical thermodynamic regime (Assumption 3), the amortized energy cost per query satisfies:
$\bar{E}(\Pi)\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}\cdot k_{B}T\ln 2-O(k_{B}T)$ (121)
where $k_{B}$ is Boltzmann’s constant and $T$ is the operating temperature.
By Proposition 4, each derivation step requires at least $k_{B}T\ln 2$ of energy dissipation, giving:
$E_{\mathrm{compute}}(q|S_{O}\cup S)\geq D_{d}(q|S_{O}\cup S)\cdot k_{B}T\ln 2\\ =H_{derive}(q\mid S_{O}\cup S)\cdot k_{B}T$ (122)
By Proposition 5, maintaining $|S|$ bits for time $t$ requires:
$E_{\mathrm{storage}}(S,t)\geq|S|\cdot k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}}$ (123)
Under the assumption that the system operates in the standard refresh regime, where the average number of refresh cycles per query satisfies $\frac{t}{N\cdot t_{\mathrm{refresh}}}=\Theta(1)$, we have:
$\frac{E_{\mathrm{storage}}(S,t)}{N}=\Theta\left(\frac{|S|}{N}\right)\cdot k_{B}T\ln 2$ (124)
Combining these bounds:
$\bar{E}(\Pi)$
$=\frac{E_{\mathrm{storage}}(S,t)}{N}+\mathbb{E}\left[E_{\mathrm{compute}}(Q|S_{O}\cup S)\right]$
$\geq\Theta(1)\cdot\left(\frac{|S|}{N}+\mathbb{E}\left[H_{\mathrm{derive}}(Q|S_{O}\cup S)\right]\right)\cdot k_{B}T\ln 2$
$=\Theta(1)\cdot\bar{I}(\Pi)\cdot k_{B}T\ln 2$
$\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}\cdot k_{B}T\ln 2-O(k_{B}T)$ (125)
where the last inequality applies Theorem 6. ∎
Corollary 7 (Per-Nit Energy Cost).
The minimum energy required to process one nit (natural unit) of semantic information content is:
$\frac{\bar{E}^{*}}{H(Q|S_{O})}=\Theta\left(\frac{k_{B}T}{L^{*}}\right)$ (126)
where $L^{*}=\Theta(\log|\mathrm{Atom}(S_{O})|)$ under the polynomial-depth regime (Assumption 4).
This follows directly from Corollary 6 and Remark 13 by dividing both sides by $H(Q|S_{O})$. ∎
Theorem 6 and its corollaries establish several fundamental insights:
The Logarithmic Efficiency Factor.
The denominator $L_{S}=\log(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}^{\sup}(P_{Q}|S_{O}\cup S))$ represents a compression factor arising from the logical structure of the knowledge base. This factor quantifies the efficiency gain from exploiting logical relationships rather than storing or computing information independently. In large knowledge systems where $|\mathrm{Atom}(S_{O})|=2^{\Omega(n)}$, this factor can be substantial, yielding energy costs that scale as $O(n)$ rather than $O(2^{n})$.
Comparison with Landauer’s Principle.
The classical Landauer bound states that erasing one bit requires at least $k_{B}T\ln 2$ of energy. Corollary 7 shows that processing one nit of semantic information requires $\Theta(k_{B}T/\log|\mathrm{Atom}|)$ of energy. The additional $1/\log|\mathrm{Atom}|$ factor reflects the efficiency gained through logical inference: derivation compresses information processing by exploiting the structure of the knowledge base.
The Storage-Computation Boundary.
The lower bound (IV-B) treats storage and computation symmetrically through their information-theoretic costs. This symmetry breaks down in the energy interpretation (Corollary 6) only through the refresh rate parameter $t_{\mathrm{refresh}}$, which is a technology-dependent quantity. The information-theoretic formulation thus captures the fundamental tradeoff independent of implementation details.
Regime Conditions.
The assumption $N=\Omega(L_{S})$ ensures that the system processes enough queries to amortize storage costs. When $N=o(L_{S})$, storage costs dominate and the bound must be modified. This corresponds to the intuition that caching is only beneficial when queries are sufficiently frequent relative to the system’s logical complexity.